# Adapting SAM via Cross-Entropy Masking for Class Imbalance in Remote Sensing Change Detection

### v1

@@ III Experimentation and Results
SAM + STFE + D 95.87 92.32
SAM + STFE + D + MSDF 95.92 92.38
SAM + STFE + D + MSDF + CEM 95.98 92.51
TABLE V: Ablation study using different masking probability values on the Levir-CD dataset, results are in %, and $\delta$ is the masking ratio.
Metric $\delta=0.2$ $\delta=0.3$ $\delta=0.4$ $\delta=0.5$ $\delta=0.6$
$mF_{1}$ 95.93 95.98 95.94 95.86 95.80
mIoUmIoU 92.41 92.51 92.44 92.29 92.19
TABLE VI: Ablation study using different losses for class imbalance on the Levir-CD dataset, results are in %.
@@
BCE (=0.7=0.7) + Dice (=0.3=0.3) 95.74 92.08
Focal ($\alpha=0.5$, $\gamma=2.0$) 95.78 92.16
Weighted-BCE ($w_{0}=0.7$, $w_{1}=1.0$) 95.87 92.31
Weighted-BCE ($w_{0}=0.3$, $w_{1}=0.7$) 95.66 91.95
CEM ($\delta=0.3$) 95.98 92.51
@@
TABLE VII: Ablation study using different pre-trained models in our methodology on the Levir-CD dataset with and without CEM, results are in %,
Method $mF_{1}$ mIoUmIoU $mF_{1}$ (CEM) mIoUmIoU (CEM)
VGG-16 [9] 95.10 91.00 95.33 91.38
ResNet-101 [8] 95.08 90.96 95.31 91.34
ResNet-152 [8] 95.09 90.98 95.39 91.48
WideResNet-101-2 [19] 95.00 90.83 95.01 90.84
Ours 95.92 92.38 95.98 92.51
III-A Implementation Details
We used PyTorch for the model training. We trained the model for 35 epochs using a stochastic gradient descent (SGD) optimizer. We use the FastSAM encoder and fine-tune its pre-trained weights. The learning rate (lr) is 0.01 for Levir-CD and CLCD, and 0.001 for WHU-CD and S2Looking. We decay the learning rate in each iteration $lr*(1-iterations/50)^{decay}$, where decay is 2.0 for Levir-CD and CLCD, and 3.0 for WHU-CD and S2Looking. Image flipping and random cropping augmentations are applied during training and images are flipped 4 times for inference to produce stable results. More details can be seen in our GitHub repo: https://github.com/humza909/SAM-CEM-CD.
III-B Change Detection Datasets
III-B1 Levir-CD [20]
This dataset contains 637 image pairs with 31,333 annotated change objects. The data is collected from Google Earth for six Texas cities between 2012 and 2016. It has 445 training, 64 validation, and 128 testing images.
III-B2 WHU-CD [21]
This is an aerial imaging dataset captured over Christchurch, New Zealand, after the 2011 earthquake, between 2012-2016. We used the dataset split of 6096 training, 762 validation, and testing images provided by [11].
III-B3 CLCD [22]
The cropland change detection dataset is captured in Guangdong, China between 2017 to 2019. It consists of 600 image pairs collected by the Gaofen-2 satellite with a split of 320 training, 120 validation and testing image pairs.
III-B4 S2Looking [23]
This is a large-scale change detection dataset with 5000 image pairs collected from satellites such as Gaofen, SuperView, and BeiJing-2. The dataset offers various challenges like misalignment and varying poses.
III-C Evaluation Metrics
To evaluate the performance of our model, we employ widely used metrics for change detection, including precision (Prec), recall (Rec), $F_{1}$ score, overall accuracy (OA), and intersection over union (IoU). The formulas are the same as given in [11].
III-D Quantitative Results
Our method, SAM-CEM-CD, achieves superior performance than other SOTA methods in the literature. For the correctness of the results, we ran our experiments three times for each dataset and averaged the results. The results for other methods in Tables III, and III are collected from [11]. Table I shows an increase of 0.48% and 0.83% in mF1 and mIoU, respectively, for the Levir-CD dataset. Similarly, we achieve 97.23% mF1 and 94.73% mIoU for the WHU-CD dataset, compared to the second-best CGNET [6], which yields 97.07% mF1 and 94.44% mIoU. We recalculated SAM-CD results for WHU-CD as given in Table I. Table II shows an approximately 1% improvement in mF1 and mIoU for CLCD against other methods, whereas for S2Looking we achieve 2.5% better performance than the SOTA methods in Table III.
III-E Visualization
We visually compare the change maps of SAM-CD and SAM-CEM-CD in Figure 3. The blue predictions are FNs, and the red predictions are FPs. We can see the reduced FNs and FPs by SAM-CEM-MCD in the figure, which shows the significance of our method over other SOTA methods.
III-F Ablation Study
III-F1 Cross-Entropy Masking Ratio
@@
Cross-entropy masking balances the impact of a highly imbalanced class ratio in change detection datasets. However, finding a reasonable masking ratio is important for optimal network performance. Higher values of masking can lead to removing too much information for the network to learn, while low values can lead to cross-entropy masking performing equivalent to cross-entropy. We show an analysis of different masking ratios in Table V and found that masking probability $\delta$ equal to 0.3 is optimal.
III-F2 Architectural Modules
We compared the impact of various modules in our architecture. We show the improvement in performance achieved due to STFE, MSDF, and CEM in Table IV, where the final architecture with all the modules performs better than all the other configurations.
III-F3 Losses for Class Imbalance
@@
We compared cross-entropy masking (CEM) with other losses widely used for class imbalance, including focal loss [24], a combination of binary cross-entropy and dice loss [25], and weighted binary cross-entropy [25] in Table VI. Compared with these, CEM performs the best, whereas weighted-BCE is the second best.
III-F4 Pre-trained Model Comparison
A comparison between pre-trained models for non-dense prediction tasks [9, 8, 19] against dense prediction tasks [14, 15] is shown in Table VII. We replaced the encoder in Figure 2 with encoders of various pre-trained models. We can see the model trained from image segmentation outperforms others.
IV Conclusion
@@
This paper presents employing the dense prediction SAM foundational model for change detection. We proposed spatial-temporal feature enhancement (STFE) and multi-scale decoder fusion (MSDF) architectural modules to enhance the detection performance. Our novel cross-entropy masking (CEM) loss achieved large gains over other losses suggested for class imbalance. Overall, we showed using an encoder from pre-trained models for dense prediction tasks along with CEM is an effective architectural design that outperforms SOTA methods.
Acknowledgement

### v2

@@ 4.6 Ablation Study
SAM + STFE + D 95.83 92.24
SAM + STFE + D + MSDFA 95.88 92.33
SAM + STFE + D + MSDFA + CEM 96.01 92.56
Architectural Modules: We compared the impact of various modules in our architecture. We demonstrate the improvement in performance over the baseline by adding MSDFA in Table 4, which fuses multi-scale information from the reconstructed change maps. The mF1 increases from 95.83% to 95.88%, and mIoU increases from 92.24% to 92.33% by using MSDFA. Training the network with CEM loss improves performance and generalizability, increasing mF1 from 95.88% to 96.01%, and mIoU from 92.33% to 92.56%.
Table 5: Ablation study on using different receptive field sizes in the model, results are for Levir-CD dataset.
@@
Receptive Field = 1 95.94 92.43
Receptive Field = 3 95.92 92.39
Receptive Field = 5 95.96 92.46
Ensemble of Receptive Field = 1, 3, 5 96.01 92.56
Ensemble of Receptive Fields: We analyze the effect of using an ensemble of different receptive field sizes. The analysis in Table 5 shows that using a combination of receptive fields $1\times 1$, $3\times 3$, and $5\times 5$ is better at capturing local and global semantics than using a single branch of one of these receptive fields. All of these experiments are performed using CEM loss.
Table 6: Ablation study on using different kernel sizes in MSDFA for different branches in an ensemble, results are for Levir-CD dataset. $k$ is the kernel size in MSDFA, and $i$ is the receptive field size in the ensemble.
Method $mF_{1}$ mIoUmIoU
MSDFA (k=3, i=1,3,5) 95.99 92.52
MSDFA (k=5, i=1,3,5) 95.99 92.53
MSDFA (k=7, i=1,3,5) 95.96 92.48
MSDFA (k=3,5,7, i=1,3,5) 96.01 92.56
Multi-Scale Decoder Fusion with Attention: For each ensemble branch in our architecture, we propose using one size larger kernel sizes in MSDFA for both channel and spatial attention. This allows the model to capture more details in the fused feature maps. We analyze this effect in Table 6, where we kept the same kernel sizes for all branches in the ensemble compared to a different kernel size for each branch. We found that employing a larger kernel $k=3,5,7$ size than the receptive field size of the ensemble branch $i=1,3,5$ performs better with 96.01% mF1 and 92.56% mIoU. The kernel size of 5 for all the ensemble branches achieves 95.99% mF1 and 92.52% mIoU, the second-best performance among other combinations.
@@
Table 7: Ablation study on using different masking probability values on the Levir-CD dataset, results are in %, and $\delta$ is the masking ratio.
Metric $\delta=0.1$ $\delta=0.2$ $\delta=0.3$ $\delta=0.4$ $\delta=0.5$ $\delta=0.6$
$mF_{1}$ 95.94 95.97 96.01 95.97 95.87 95.85
mIoUmIoU 92.43 92.47 92.56 92.49 92.32 92.27
@@
Cross-Entropy Masking Ratio: Cross-entropy masking balances the impact of a highly imbalanced class ratio in change detection datasets. However, finding a reasonable masking ratio is important for optimal network performance. Higher values of masking can lead to removing too much information for the network to learn, while low values can lead to cross-entropy masking performing equivalent to cross-entropy. We show an analysis of different masking ratios in Table 7 and found that masking probability $\delta$ equal to 0.3 is optimal.
Table 8: Ablation study using different losses for class imbalance on the Levir-CD dataset, results are in %.
Method $mF_{1}$ mIoUmIoU
BCE (=0.7=0.7) + Dice (=0.3=0.3) 95.81 92.21
Focal ($\alpha=0.5$, $\gamma=2.0$) 95.46 91.61
Weighted-BCE ($w_{0}=1.0$, $w_{1}=20.0$) 92.64 86.97
Weighted-BCE ($w_{0}=0.3$, $w_{1}=0.7$) 95.75 92.10
CEM ($\delta=0.3$) 96.01 92.56
@@
Losses for Class Imbalance: We compared cross-entropy masking (CEM) with other losses widely used for class imbalance, including focal loss [15], a combination of binary cross-entropy (BCE) and dice loss [10], and weighted binary cross-entropy [10] in Table 8. These loss functions adjust the weights of the minority class to penalize it more for false detections. This results in overfitting and even worse performance, as shown in weighted-BCE with equal to 1.0 and $w_{1}$ equal to 20.0, accounting for the exact class-imbalance in Levir-CD dataset. In weighted-BCE, $w_{0}$ is the unchanged class and $w_{1}$ is the changed class. Compared to these, CEM outperforms BCE and other loss functions for class-imbalance. A combination of BCE and dice loss is the second best with 95.81% mF1 and 92.21% mIoU.
5 Conclusion
@@
This paper presents a new architecture to learn change patterns at different scales by using an ensemble of multi-receptive fields. We propose an ensemble of spatial-temporal feature enhancement (STFE), decoder, and multi-scale decoder fusion with attention (MSDFA) to effectively capture local and global contextual change patterns in bi-temporal images. The proposed novel cross-entropy masking (CEM) loss achieved large gains over other losses suggested for class imbalance. Our method demonstrated significant performance gains that outperformed recent remote sensing change detection (RSCD) state-of-the-art (SOTA) methods.
Acknowledgment